# Semantic Agreement Enables Efficient Open-Ended LLM Cascades

### v1

@@ 1 Introduction
Large language models (LLMs) have enabled impressive progress across a range of language tasks; however, this progress comes at a steep computational cost. Larger models typically produce higher-quality outputs but are slower, more expensive, and less scalable for real-time or large-scale use. To mitigate this cost-quality tradeoff, cascade systems have emerged as a practical deployment strategy: route inputs to smaller models whenever possible, and defer to larger models only when necessary.  (Chen et al., 2024; Yue et al., 2024; Kolawole et al., 2025)
@@
To address these challenges, we introduce a simple, training-free alternative: using semantic agreement among multiple model outputs as a deferral signal. When independently generated outputs are semantically consistent—even if lexically different—their agreement suggests the underlying meaning is reliable. In contrast, semantic divergence signals uncertainty, suggesting that deferring to a larger model may be warranted. Our method requires no model-specific tuning, no access to internals, and generalizes across model families and API versions—while capturing aspects of output reliability that token-level confidence methods miss entirely.
@@ 3.1 Deferral Protocol
A semantic cascade comprises $n$ lightweight ensemble models, $M_{1},\dots,M_{n}$, and a larger, high-capacity target model $M_{T}$. Given an input $x$, each ensemble model produces a full response $y^{(i)}=M_{i}(x)$. A deferral score function $s(y^{(i)},~\dots,~y^{(n)})$ then determines whether to defer to $M_{T}$, based on semantic agreement across outputs. We adopt the convention that higher scores reflect greater certainty; the system retains ensemble predictions for high scores and defers for low scores. When not deferring, the system selects a final output $y=y^{(i)}$ corresponding to the response with highest score according to an output scoring function $o(y^{(i)},~y^{(1)},~\dots,~y^{(n)})$.
@@ 3.2 Semantic Agreement Signals
In an ensemble setting, we use these similarity metrics as the output scoring function $o$ by computing the mean pairwise similarity between $y^{(i)}$ and all other $y^{(j)}$. We then use the output scores to determine the deferral score by $s=\max_{i}o(y^{(i)},~y^{(1)},~\dots,~y^{(n)})$. This approach identifies the output with highest agreement across the ensemble, while providing a confidence signal for deferral decisions.
@@ 5.3 Mechanistic Advantages of Semantics
Semantic agreement’s strength stems from two areas. First, semantic similarity’s ability to reliably select a strong response from the ensemble’s several outputs leads to high baseline performances, an effect which is especially apparent in top performing ensembles (Figure˜1). Second, semantic similarity provides a stronger indication of when to defer. Even when baseline ensemble performance is not higher than a token-level cascade’s baseline—which occurs when the semantic ensemble is comprised of unilaterally weaker (and often cheaper) models—semantic similarity can still outperform token-level methods through more efficient deferral choices arising from accurate identification of (un)reliable model outputs (Figure˜2).
@@ Appendix B BLEURT Model Sizes
There are several different official BLEURT model sizes available. We use BLEURT-20 in our experiments, which has 30 layers and 579M parameters, but there are also three lossily compressed versions available: BLEURT-20-D12 (12 layers, 167M parameters), BLEURT-20-D6 (6 layers, 45M parameters), and BLEURT-20-D3 (3 layers, 30M parameters). We find that on domains with longer output lengths, the size (quality) of the BLEURT model becomes important, but is largely irrelevant at very small scales (Figure˜3).
@@ Appendix E Difficulties with Assessing Summarization
Further, we note that on XLSum no deferral strategy significantly improves on random deferral—especially at smaller scales—leading to flat deferral curves and baseline performance becoming the dominating factor (Figure˜5).
@@ Appendix F Ensembling Token-level Models
Avg Sum Avg Sum Avg Sum Avg Sum
Llama3.1-8B -0.2378 -1.3165 -0.1209 -0.6837 -0.1162 -2.9247 -0.2189 -21.2436
Llama3.2-3B -0.2990 -1.2517 -0.1467 -0.7346 -0.1405 -3.5432 -0.2455 -29.0617
Llama3.2-1B -0.6003 -2.6633 -0.2863 -1.1109 -0.2442 -5.9975 -0.3268 -35.6698
Qwen2.5-7B -0.1929 -0.8508 -0.0288 -0.1835 -0.0461 -1.2387 -0.1945 -11.7805
Qwen2.5-3B -0.2733 -1.1296 -0.0541 -0.3209 -0.0664 -1.7102 -0.3397 -27.4709
Qwen2.5-1.5B -0.3850 -1.7507 -0.0882 -0.5509 -0.1366 -3.5275 -0.3694 -15.1451
Qwen2.5-0.5B -0.6137 -3.7378 -0.1955 -1.1930 -0.2966 -7.5582 -0.4268 -21.1389
Gemma3-1B -0.2602 -1.0226 -0.0667 -0.3466 -0.0635 -1.5018 -0.1625 -12.5079
FLAN-T5-Large -0.8618 -4.7093 -0.1070 -0.9355 -0.4093 -10.6007 -0.3233 -20.2584
mT0-Large -1.2154 -5.3315 -0.1285 -0.5957 -0.4325 -12.2072 -0.4301 -24.4393
@@ Appendix G TriviaQA: Performance Gap and Oracle Insights
It is tempting to assume the performance gap between 3B parameter models and ensembles of 1B parameter models arises because the 3B parameter models simply encode more information than the smaller ensemble models. Surprisingly, however, the oracle deferral curves for this task suggest that this explanation is incomplete and that other dynamics are at play (Figure˜6). We hypothesize that the gap in performance originates not simply from a lack of encoded information, but also due to the difficulty of the ensemble deferral task, which requires not only identifying when to defer to a large model, but also selecting the best answer from the ensemble’s multiple outputs when not deferring. This challenge is particularly pronounced in TriviaQA, where answers are short and baseline ensemble models perform poorly—leading to frequent disagreement from which semantic similarity struggles to extract meaningful signal.
@@ H.2 AUC-DF Values
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.6928 0.6996 0.7032 0.7023 0.1
Qwen2.5-3B 0.669 0.6766 0.683 0.6826 0.3
Qwen2.5-1.5B 0.6525 0.6625 0.6701 0.6688 0.2
Qwen2.5-0.5B 0.5624 0.5766 0.5842 0.583 0.5
Gemma3-1B 0.6465 0.6573 0.6646 0.6641 0.3
Llama3.2-1B 0.5988 0.6149 0.6217 0.6206 0.3
Llama3.2-3B 0.6718 0.6797 0.6857 0.6854 0.2
Llama3.1-8B 0.7031 0.7065 0.7093 0.7089 0.2
FLAN-T5-Large 0.5611 0.5857 0.5886 0.5873 0.2
mT0-Large 0.5863 0.6044 0.6098 0.6085 0.3
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.6702 0.6738 0.6729 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6616 0.6613 0.6627 0.2
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6595 0.661 0.6606 0.2
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6552 0.6532 0.6538 0.1
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6486 0.6445 0.6484 1.0
Llama3.1-8B, Qwen2.5-7B 0.7075 0.7096 0.7084 0.1
Llama3.1-8B, Qwen2.5-3B 0.6976 0.7002 0.7025 0.6
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6801 0.6758 0.6758 0.2
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6808 0.6839 0.7019 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.6621 0.6641 0.6637 0.6654 0.6806 0.6686
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.6632 0.6664 0.6662 0.6667 0.6861 0.6705
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.6626 0.6645 0.6648 0.6651 0.6854 0.6717
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.6493 0.6521 0.6499 0.6524 0.6747 0.6581
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.6618 0.6637 0.6642 0.6657 0.6877 0.6702
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.6609 0.6637 0.6636 0.6649 0.6889 0.6711
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.7046 0.7048 0.7052 0.7052 0.7141 0.7063
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.7045 0.7054 0.7051 0.7048 0.716 0.7059
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.7043 0.705 0.7052 0.705 0.7141 0.7064
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.6835 0.6849 0.6848 0.6857 0.7005 0.6892
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.7438 0.7454 0.7453 0.7454 0.0
Qwen2.5-3B 0.7382 0.739 0.7397 0.7397 0.1
Qwen2.5-1.5B 0.736 0.7383 0.7391 0.739 0.1
Qwen2.5-0.5B 0.7152 0.7193 0.7224 0.7216 0.1
Gemma3-1B 0.7278 0.7299 0.7316 0.7321 0.5
Llama3.2-1B 0.7114 0.715 0.7178 0.7176 0.6
Llama3.2-3B 0.7379 0.7395 0.7404 0.7402 0.1
Llama3.1-8B 0.7433 0.7439 0.7445 0.7445 0.1
FLAN-T5-Large 0.7174 0.7221 0.7268 0.7268 0.1
mT0-Large 0.7051 0.7102 0.7149 0.7143 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.7354 0.7359 0.7377 0.7
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7303 0.7302 0.7306 0.0
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.7301 0.73 0.731 0.2
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.7348 0.7348 0.7348 0.2
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7277 0.7261 0.7269 0.0
Llama3.1-8B, Qwen2.5-7B 0.7453 0.7452 0.7453 0.1
Llama3.1-8B, Qwen2.5-3B 0.7429 0.7431 0.7436 0.7
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.7425 0.7402 0.7414 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7382 0.7364 0.7433 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.7362 0.7362 0.7358 0.7364 0.7397 0.737
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.737 0.7372 0.7359 0.7373 0.7408 0.7379
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.7385 0.7382 0.7378 0.7387 0.742 0.7395
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.7342 0.7346 0.7345 0.7347 0.7384 0.7359
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.7368 0.7368 0.7364 0.737 0.7406 0.7373
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.7374 0.7375 0.7362 0.7374 0.7418 0.7388
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.7454 0.7453 0.7448 0.7452 0.747 0.7458
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.7448 0.7445 0.7442 0.7445 0.7464 0.7451
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.7451 0.7449 0.7451 0.7453 0.7461 0.7454
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.7403 0.7404 0.7399 0.7403 0.7431 0.7415
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Qwen2.5-7B 0.6821 0.6897 0.6931 0.6921 0.1
Qwen2.5-3B 0.6676 0.6728 0.6784 0.6775 0.2
Qwen2.5-1.5B 0.6592 0.6669 0.672 0.6711 0.3
Qwen2.5-0.5B 0.5973 0.6116 0.6173 0.6164 0.5
Gemma3-1B 0.6654 0.6724 0.6754 0.6744 0.1
Llama3.2-1B 0.642 0.6549 0.6601 0.6592 0.2
Llama3.2-3B 0.6777 0.6833 0.6879 0.6873 0.2
Llama3.1-8B 0.693 0.6953 0.6979 0.698 0.6
FLAN-T5-Large 0.5802 0.6012 0.605 0.6039 0.1
mT0-Large 0.6095 0.6259 0.6305 0.6287 0.4
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile
Gemma3-1B, Qwen2.5-1.5B 0.6772 0.6774 0.677 0.1
Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6766 0.6763 0.6764 0.1
Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6697 0.6688 0.6672 0.1
Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6619 0.6625 0.6594 0.1
Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6639 0.6644 0.6624 0.2
Llama3.1-8B, Qwen2.5-7B 0.6962 0.6982 0.6976 0.7
Llama3.1-8B, Qwen2.5-3B 0.6885 0.6911 0.6935 1.0
Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6741 0.6733 0.6729 0.1
Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6825 0.6831 0.6936 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B
0.6724 0.672 0.6725 0.6727 0.6835 0.6749
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large
0.6712 0.6714 0.6709 0.6718 0.6861 0.6733
Qwen2.5-1.5B, Gemma3-1B,
FLAN-T5-Large
0.6673 0.6695 0.6688 0.6688 0.683 0.6717
Qwen2.5-1.5B, Llama3.2-1B,
mT0-Large
0.6649 0.6635 0.6648 0.6651 0.6787 0.6664
Gemma3-1B, Qwen2.5-1.5B,
Llama3.2-1B, mT0-Large
0.6736 0.6727 0.6737 0.6739 0.6878 0.6752
Gemma3-1B, Qwen2.5-1.5B,
mT0-Large, FLAN-T5-Large
0.6678 0.6705 0.6696 0.6706 0.6864 0.6709
Llama3.1-8B, Qwen2.5-7B,
FLAN-T5-Large
0.6938 0.6933 0.6939 0.6938 0.7003 0.6948
Llama3.1-8B, Qwen2.5-7B,
mT0-Large
0.6944 0.6931 0.6936 0.6932 0.7013 0.6938
Llama3.1-8B, Qwen2.5-7B,
Qwen2.5-1.5B
0.6929 0.6933 0.6934 0.6933 0.6996 0.6938
Llama3.2-3B, Qwen2.5-3B,
Gemma3-1B
0.6859 0.6865 0.6861 0.6867 0.6961 0.6887

### v3

@@ 1 Introduction
Large language models (LLMs) have enabled impressive progress across a range of language tasks; however, this progress comes at a steep computational cost. Larger models typically produce higher-quality outputs but are slower, more expensive, and less scalable for real-time or large-scale use. To mitigate this cost-quality tradeoff, cascade systems have emerged as a practical deployment strategy: route inputs to smaller models whenever possible, and defer to larger models only when necessary  (Chen et al., 2024; Yue et al., 2024; Kolawole et al., 2025).
Figure 1: Overview of the semantic cascade framework. An ensemble of small models generates responses to a user query. Semantic similarity metrics assess agreement among outputs. High semantic agreement indicates reliability: the system returns the highest-scoring ensemble response. Low agreement signals uncertainty: the query defers to a larger target model, ensuring that expensive model is called only when necessary.
@@
To address these challenges, we introduce a simple, training-free alternative: using semantic agreement among multiple model outputs as a deferral signal (overview in Figure 1). When independently generated outputs are semantically consistent—even if lexically different—their agreement suggests the underlying meaning is reliable. In contrast, semantic divergence signals uncertainty, suggesting that deferring to a larger model may be warranted. Our method requires no model-specific tuning, no access to internals, and generalizes across model families and API versions—while capturing aspects of output reliability that token-level confidence methods miss entirely.
@@ 3.1 Deferral Protocol
A semantic cascade comprises $n$ lightweight ensemble models, $M_{1},\dots,M_{n}$, and a larger, high-capacity target model $M_{T}$. Given an input $x$, each ensemble model produces a full response $y^{(i)}=M_{i}(x)$. A deferral score function $s\left(y^{(i)},~\dots,~y^{(n)}\right)$ then determines whether to defer to $M_{T}$, based on semantic agreement across outputs. We adopt the convention that higher scores reflect greater certainty; the system retains ensemble predictions for high scores and defers for low scores. When not deferring, the system selects a final output $y=y^{(i)}$ corresponding to the response with highest score according to an output scoring function $o\left(y^{(i)},~y^{(1)},~\dots,~y^{(n)}\right)$.
@@ 3.2 Semantic Agreement Signals
In an ensemble setting, we use these similarity metrics as the output scoring function $o$ by computing the mean pairwise similarity between $y^{(i)}$ and all other $y^{(j)}$. We then use the output scores to determine the deferral score by $s=\max_{i}o\left(y^{(i)},~y^{(1)},~\dots,~y^{(n)}\right)$. This approach identifies the output with highest agreement across the ensemble, while providing a confidence signal for deferral decisions.
@@ 5.3 Mechanistic Advantages of Semantics
Semantic agreement’s strength stems from two areas. First, semantic similarity’s ability to reliably select a strong response from the ensemble’s several outputs leads to high baseline performances, an effect which is especially apparent in top performing ensembles (Figure˜2). Second, semantic similarity provides a stronger indication of when to defer. Even when baseline ensemble performance is not higher than a token-level cascade’s baseline—which occurs when the semantic ensemble is comprised of unilaterally weaker (and often cheaper) models—semantic similarity can still outperform token-level methods through more efficient deferral choices arising from accurate identification of (un)reliable model outputs (Figure˜3).
@@ Appendix B BLEURT Model Sizes
There are several different official BLEURT model sizes available. We use BLEURT-20 in our experiments, which has 30 layers and 579M parameters, but there are also three lossily compressed versions available: BLEURT-20-D12 (12 layers, 167M parameters), BLEURT-20-D6 (6 layers, 45M parameters), and BLEURT-20-D3 (3 layers, 30M parameters). We find that on domains with longer output lengths, the size (quality) of the BLEURT model becomes important, but is largely irrelevant at very small scales (Figure˜4).
@@ Appendix E Difficulties with Assessing Summarization
Further, we note that on XLSum no deferral strategy significantly improves on random deferral—especially at smaller scales—leading to flat deferral curves and baseline performance becoming the dominating factor (Figure˜6).
@@ Appendix G TriviaQA: Performance Gap and Oracle Insights
It is tempting to assume the performance gap between 3B parameter models and ensembles of 1B parameter models arises because the 3B parameter models simply encode more information than the smaller ensemble models. Surprisingly, however, the oracle deferral curves for this task suggest that this explanation is incomplete and that other dynamics are at play (Figure˜7). We hypothesize that the gap in performance originates not simply from a lack of encoded information, but also due to the difficulty of the ensemble deferral task, which requires not only identifying when to defer to a large model, but also selecting the best answer from the ensemble’s multiple outputs when not deferring. This challenge is particularly pronounced in TriviaQA, where answers are short and baseline ensemble models perform poorly—leading to frequent disagreement from which semantic similarity struggles to extract meaningful signal.
@@ H.2 AUC-DF Values
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.6928 0.6996 0.7032 0.7023 0.1 Qwen2.5-3B 0.669 0.6766 0.683 0.6826 0.3 Qwen2.5-1.5B 0.6525 0.6625 0.6701 0.6688 0.2 Qwen2.5-0.5B 0.5624 0.5766 0.5842 0.583 0.5 Gemma3-1B 0.6465 0.6573 0.6646 0.6641 0.3 Llama3.2-1B 0.5988 0.6149 0.6217 0.6206 0.3 Llama3.2-3B 0.6718 0.6797 0.6857 0.6854 0.2 Llama3.1-8B 0.7031 0.7065 0.7093 0.7089 0.2 FLAN-T5-Large 0.5611 0.5857 0.5886 0.5873 0.2 mT0-Large 0.5863 0.6044 0.6098 0.6085 0.3
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.6702 0.6738 0.6729 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6616 0.6613 0.6627 0.2 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6595 0.661 0.6606 0.2 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6552 0.6532 0.6538 0.1 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6486 0.6445 0.6484 1.0 Llama3.1-8B, Qwen2.5-7B 0.7075 0.7096 0.7084 0.1 Llama3.1-8B, Qwen2.5-3B 0.6976 0.7002 0.7025 0.6 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6801 0.6758 0.6758 0.2 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6808 0.6839 0.7019 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6621 0.6641 0.6637 0.6654 0.6806 0.6686 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6632 0.6664 0.6662 0.6667 0.6861 0.6705 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.6626 0.6645 0.6648 0.6651 0.6854 0.6717 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6493 0.6521 0.6499 0.6524 0.6747 0.6581 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6618 0.6637 0.6642 0.6657 0.6877 0.6702 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.6609 0.6637 0.6636 0.6649 0.6889 0.6711 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.7046 0.7048 0.7052 0.7052 0.7141 0.7063 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7045 0.7054 0.7051 0.7048 0.716 0.7059 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.7043 0.705 0.7052 0.705 0.7141 0.7064 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.6835 0.6849 0.6848 0.6857 0.7005 0.6892
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.7438 0.7454 0.7453 0.7454 0.0 Qwen2.5-3B 0.7382 0.739 0.7397 0.7397 0.1 Qwen2.5-1.5B 0.736 0.7383 0.7391 0.739 0.1 Qwen2.5-0.5B 0.7152 0.7193 0.7224 0.7216 0.1 Gemma3-1B 0.7278 0.7299 0.7316 0.7321 0.5 Llama3.2-1B 0.7114 0.715 0.7178 0.7176 0.6 Llama3.2-3B 0.7379 0.7395 0.7404 0.7402 0.1 Llama3.1-8B 0.7433 0.7439 0.7445 0.7445 0.1 FLAN-T5-Large 0.7174 0.7221 0.7268 0.7268 0.1 mT0-Large 0.7051 0.7102 0.7149 0.7143 0.1
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.7354 0.7359 0.7377 0.7 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7303 0.7302 0.7306 0.0 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.7301 0.73 0.731 0.2 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.7348 0.7348 0.7348 0.2 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7277 0.7261 0.7269 0.0 Llama3.1-8B, Qwen2.5-7B 0.7453 0.7452 0.7453 0.1 Llama3.1-8B, Qwen2.5-3B 0.7429 0.7431 0.7436 0.7 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.7425 0.7402 0.7414 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7382 0.7364 0.7433 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.7362 0.7362 0.7358 0.7364 0.7397 0.737 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.737 0.7372 0.7359 0.7373 0.7408 0.7379 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.7385 0.7382 0.7378 0.7387 0.742 0.7395 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7342 0.7346 0.7345 0.7347 0.7384 0.7359 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.7368 0.7368 0.7364 0.737 0.7406 0.7373 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.7374 0.7375 0.7362 0.7374 0.7418 0.7388 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.7454 0.7453 0.7448 0.7452 0.747 0.7458 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.7448 0.7445 0.7442 0.7445 0.7464 0.7451 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.7451 0.7449 0.7451 0.7453 0.7461 0.7454 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.7403 0.7404 0.7399 0.7403 0.7431 0.7415
@@
Model Random Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Qwen2.5-7B 0.6821 0.6897 0.6931 0.6921 0.1 Qwen2.5-3B 0.6676 0.6728 0.6784 0.6775 0.2 Qwen2.5-1.5B 0.6592 0.6669 0.672 0.6711 0.3 Qwen2.5-0.5B 0.5973 0.6116 0.6173 0.6164 0.5 Gemma3-1B 0.6654 0.6724 0.6754 0.6744 0.1 Llama3.2-1B 0.642 0.6549 0.6601 0.6592 0.2 Llama3.2-3B 0.6777 0.6833 0.6879 0.6873 0.2 Llama3.1-8B 0.693 0.6953 0.6979 0.698 0.6 FLAN-T5-Large 0.5802 0.6012 0.605 0.6039 0.1 mT0-Large 0.6095 0.6259 0.6305 0.6287 0.4
@@
Ensemble Chow-Sum Chow-Avg Best Chow-Quantile Best Quantile Gemma3-1B, Qwen2.5-1.5B 0.6772 0.6774 0.677 0.1 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6766 0.6763 0.6764 0.1 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6697 0.6688 0.6672 0.1 Gemma3-1B, Qwen2.5-1.5B, FLAN-T5-Large 0.6619 0.6625 0.6594 0.1 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6639 0.6644 0.6624 0.2 Llama3.1-8B, Qwen2.5-7B 0.6962 0.6982 0.6976 0.7 Llama3.1-8B, Qwen2.5-3B 0.6885 0.6911 0.6935 1.0 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6741 0.6733 0.6729 0.1 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6825 0.6831 0.6936 1.0
@@
Ensemble BLEU ROUGE-1 ROUGE-2 ROUGE-L BLEURT SBERT Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B 0.6724 0.672 0.6725 0.6727 0.6835 0.6749 Gemma3-1B, Qwen2.5-1.5B, mT0-Large 0.6712 0.6714 0.6709 0.6718 0.6861 0.6733 Qwen2.5-1.5B, Gemma3-1B, FLAN-T5-Large 0.6673 0.6695 0.6688 0.6688 0.683 0.6717 Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6649 0.6635 0.6648 0.6651 0.6787 0.6664 Gemma3-1B, Qwen2.5-1.5B, Llama3.2-1B, mT0-Large 0.6736 0.6727 0.6737 0.6739 0.6878 0.6752 Gemma3-1B, Qwen2.5-1.5B, mT0-Large, FLAN-T5-Large 0.6678 0.6705 0.6696 0.6706 0.6864 0.6709 Llama3.1-8B, Qwen2.5-7B, FLAN-T5-Large 0.6938 0.6933 0.6939 0.6938 0.7003 0.6948 Llama3.1-8B, Qwen2.5-7B, mT0-Large 0.6944 0.6931 0.6936 0.6932 0.7013 0.6938 Llama3.1-8B, Qwen2.5-7B, Qwen2.5-1.5B 0.6929 0.6933 0.6934 0.6933 0.6996 0.6938 Llama3.2-3B, Qwen2.5-3B, Gemma3-1B 0.6859 0.6865 0.6861 0.6867 0.6961 0.6887